### v2

# Coupled dimerized alternating-bond quantum spin chains in the distorted honeycomb-lattice magnet Cu5SbO6

Abstract
We analyze powder-averaged inelastic neutron scattering and magnetization data for the distorted honeycomb compound Cu5SbO6 using a first-order dimer expansion calculation and quantum Monte Carlo simulations. We show that, in contrast to the previously proposed honeycomb lattice model, Cu5SbO6 accommodates interacting dimerized spin chains with alternating ferromagnetic-antiferromagnetic couplings along the chain. Moreover, unlike the typical couplings observed in other Cu2+-based distorted honeycomb magnets, the spin chains in Cu5SbO6 primarily couple through an antiferromagnetic coupling that arises between the honeycomb layers, rather than the expected interchain coupling in the layers. This finding reveals a different magnetic coupling scheme for Cu5SbO6. In addition, utilizing x-ray spectroscopy and transmission electron microscopy, we also refine the crystal structure and stacking-fault model of the compound.
I Introduction
Dimerized quantum magnets have constituted a fertile ground for exploring a plethora of exotic emergent quantum phenomena in magnetic insulators [1, 2]. The most renowned aspect of such magnets is arguably quantum criticality. One of the quantum phase transitions is achieved in the field-induced Bose-Einstein condensation of magnons realized in several compounds [3, 4, 5, 6, 7, 8, 9, 10]. Novel phenomena in the vicinity of a quantum critical point separating a dimerized disordered phase from a long-range ordered phase such as dimensional reduction [11] and longitudinal excitation modes [12] have been observed. When frustration between exchange couplings among orthogonally oriented dimers are present as in the Shastry-Sutherland magnet, several other intriguing effects have been reported including topologically protected chiral edge modes of triplet excitations [13], pressure-induced 4-spin plaquette singlet state [14, 15], field-induced spin-nematic phase [16], incomplete devil’s staircase [17] and crystallization of triplet bound states in magnetization plateaus [18]. Dimerized magnets are built upon spin dimers consisting of a pair of spins interacting with an intra-dimer antiferromagnetic (AFM) exchange coupling. The dimer, for the case of $S=\frac{1}{2}$, exhibits a non-magnetic singlet ground state and degenerate triplet excited states in zero magnetic field which are separated by a finite energy gap. Dimers that interact with each other with a ferromagnetic (FM) exchange coupling in a one-dimensional (1D) network, forming an alternating FM-AFM spin chain, form another distinct class of dimerized magnets. Magnetic materials that are characterized as quasi-1D FM-AFM dimerized spin chains include, as examples, copper nitrate [19], BaCu2V2O8 [20], CuNb2O6 [21], (CH3)2NH2CuCl3 [22], etc.
Figure 1: (a) A monoclinic unit cell of Cu5SbO6 is displayed in which spin chains consisted of SbO6 octahedra (in silver) and edge-sharing Cu2O6 double plaquettes (in green) are aligned along the $a$-axis. The unit cell boundary is indicated by the dashed lines. (b) A chain segment with alternating $J_{1}$ and $J_{2}$ couplings is shown. The dashed purple and solid yellow lines represent intra-dimer $J_{1}$ and inter-dimer $J_{2}$ couplings, respectively. (c) The abab-plane projection of a single Cu2SbO6 layer is displayed. $B_{1}$ = 2.314 Å is the Cu-O bond perpendicular to the plaquette whereas $B_{2}$ = 1.994 Å and $B_{3}$ = 2.043 Å are the Cu-O bonds within the plaquette. The dot-dash turquoise lines represent inter-chain $J_{3}$ couplings. (d) A single layer projected on to the acac-plane is shown, where the dotted orange lines represent inter-chain $J_{4}$ couplings perpendicular to the honeycomb planes. The super-exchange path responsible for the $J_{4}$ coupling in Cu5SbO6 is displayed in (e) along with the corresponding path in Na2Cu2TeO6. The crystal structure illustrations are generated by VESTA [23].
To our knowledge, the archetypes of the quasi-1D $S=\frac{1}{2}$ alternating FM-AFM dimerized spin chain realized in a distorted honeycomb lattice are Na3Cu2SbO6 [24, 25, 26, 27, 28], Li3Cu2SbO6 [29, 30], and Na2Cu2TeO6 [26, 28, 31, 32, 33]. It should be noted that, despite the similarity in their crystal structures, spin chains in Li3Cu2SbO6 are fragmented due to a considerable site interchange between magnetic Cu2+ and non-magnetic Li1+ ions [34, 35, 29, 30]. We further note that a related compound Li2Cu2TeO6 has also been synthesized [36] but its magnetic characterization still awaits studies. The crystal structure of these compounds can be viewed as a stacking of Na1+ layers alternating with Cu2MMO6 layers where $M=\mathrm{Sb},\mathrm{Te}$ [37, 38, 39, 40]. A Cu2MMO6 layer is built upon edge-sharing elongated CuO6 octahedra forming a honeycomb lattice of Cu2+ ions with $M$ ions positioned at the center of hexagons. The honeycomb lattice is distorted due to a strong Jahn-Teller effect on oxygen octahedra of Cu2+ ions. Magnetic characteristics of these archetypal compounds are extensively studied by means of experiments as well as first-principles calculations. Experimental investigations include, for instance, inelastic neutron scattering (INS) [25, 31, 32], 23Na nuclear magnetic resonance (NMR) [41, 27], and thermodynamics measurements [38, 24, 42, 28, 43]. The experimental studies reveal a finite energy gap with the non-magnetic singlet ground state in these systems. In particular, INS experiments unambiguously show that, at base temperatures, dimers are formed between third-nearest-neighbor rather than first-nearest-neighbor spins. Surprisingly, fitting the energy dispersion of the triplet excitations yields minimal AFM $J_{3}$ coupling and FM $J_{2}$ coupling that are weaker than AFM $J_{1}$ even though the bond length of $J_{1}$ is longer than that of $J_{2}$. Here, as indicated in Figs. 1(b) and 1(c), $J_{2}$ and $J_{3}$ denote the couplings between the first-nearest-neighbor spins whereas $J_{1}$ denotes the coupling between the third-nearest-neighbor spins.
First-principles calculations employing density functional theory (DFT) [38, 42, 26, 28, 33] and density-matrix renormalization group (DMRG) [33] provide a comprehensive microscopic description for such a peculiar behavior of these exchange couplings deduced from experiments. The DFT results suggest that the magnetism of these compounds is primarily governed by the half-filled Cu2+ $3d_{x^{2}-y^{2}}$ orbital, where the local $x$- and $y$-axis almost align along Cu2+-O2- bonds in the equatorial plane of the octahedron. This explains naturally the subordinate magnitude of $J_{3}$ coupling since the Cu2+-O2--Cu2+ path for $J_{3}$ incorporates a Cu2+-O2- bond (bond $B_{1}$ in Fig. 1(c)) that is elongated and perpendicular to the Cu2+ $3d_{x^{2}-y^{2}}$ orbital. That the magnitude of $J_{1}$ coupling is much greater than that of $J_{2}$ coupling owes to the direct overlap of Cu2+ $3d_{x^{2}-y^{2}}$ orbitals with O2- $2p$ orbitals along the Cu2+-O2--…-O2--Cu2+ path across the MMO6 octahedron. On the other hand, as displayed in Fig. 1(b), since the angle of the Cu2+-O2--Cu2+ path for $J_{2}$ is close to 90, a pair of O2- $2p$ orbitals is almost orthogonal resulting in a weaker $J_{2}$. In accordance with Goodenough-Kanamori-Anderson rules [44, 45, 46, 47], the sign of $J_{2}$ coupling is expected to be FM.
Another compound in the family for exploring physics of the quasi-1D alternating dimerized spin chains has been reported in Ref. [48]. The authors of Ref. [48] have synthesized a Delafossite-derived compound Cu5SbO6 ($^{1+}_{3}Cu$$^{2+}_{2}$SbO6) that realizes a distorted honeycomb lattice of Cu2+ spins. The crystal structure of Cu5SbO6 is built from a monoclinic unit cell, as shown in Fig. 1(a), with the space group $C2/c$ (no. 15) and the lattice parameters $a$ = 8.923 Å, $b$ = 5.593 Å, $c$ = 11.845 Å, and $\beta$ = 103.585. The structure can also be viewed as a stacking of magnetic Cu2SbO6 layers alternating with non-magnetic Cu1+ layers. According to the measurements performed in Ref. [48], powder sample of Cu5SbO6 develops stacking faults which manifest as diffusive streaks along the $c^{*}$-axis in a selected-area electron diffraction pattern. The stacking faults can be effectively improved by elevating synthesis temperature. The temperature dependence of magnetic susceptibility and magnetic heat capacity indicates a formation of spin dimers as temperature is lowered below 120 K. An approximated energy gap of $\sim$190 K is obtained by fitting susceptibility data with an isolated dimer model. While it is plausible to assume a similar $J_{1}$-$J_{2}$-$J_{3}$ magnetic coupling scheme for Cu5SbO6, a DFT calculation [49] suggests a slightly different model. First of all, DFT calculation confirms the alternation of AFM $J_{1}$ and FM $J_{2}$ as in the case for related compounds. However, it is found that the AFM inter-plane coupling $J_{4}$ (see Fig. 1(d)) is significant, and $J_{3}$ is even smaller than in the related compounds. By inspecting the crystal structures of Cu5SbO6 and Na2Cu2TeO6 in between the honeycomb planes, as shown in Fig. 1(e), one can see that the two angles Cu2+-O2--Cu1+ of the super-exchange path Cu2+-O2--Cu1+-O2--Cu2+ in Cu5SbO6 amount to about 120 which facilitates the electron exchange more efficiently between the two Cu2+ ions compared to Na2Cu2TeO6 where the angles Cu2+-O2--Na1+ are close to 90.
In this article, using neutron spectroscopy and magnetometry we set out to conclude that the $J_{1}$-$J_{2}$-$J_{4}$ model is more appropriate for Cu5SbO6. The content of this article is organized as follows. We first present studies of the crystal structure and stacking faults of the compound employing x-ray diffraction, x-ray absorption fine structure, and transmission electron microscopy techniques. Based upon the exchange coupling parameters determined from neutron data, quantum Monte Carlo simulations are performed to compute model temperature-dependent susceptibility and field-dependent magnetization which are, in turn, used to verify with experimental data. Finally, we present an analysis of the powder-averaged inelastic neutron scattering data utilizing a first-order dimer expansion calculation for both $J_{1}$-$J_{2}$-$J_{3}$ and $J_{1}$-$J_{2}$-$J_{4}$ models.
II Experimental details
II.1 Crystal growth and characterization
Powder samples of Cu5SbO6 were prepared by a conventional solid-state reaction from a stoichiometric mixture of CuO and Sb2O5. A homogeneous mixture of the precursors was placed in an alumina crucible and heated in air at 950 C for 24 h followed by an intermediate grinding and re-heating at 1000 C for another 24 h. Single-crystal samples were grown from the obtained powder samples using the vertical-gradient freezing technique. The powder sample with mass $\sim$10 g was loaded into a pointed-bottom zirconia crucible and sealed with cement to prevent sample leakage during the growth process. We note that alumina crucibles do not work well with Cu5SbO6 due to sample penetration during crystallization from a liquid phase. The sealed crucible placed inside a quartz tube was positioned in a furnace where the temperature was $\sim$1160 C (the melting point) and uniform across the sample inside the crucible. The furnace was set up so that the temperature gradient below the starting point of the crucible was about -10 C/cm. The heating sequence started with increasing the temperature with a rate of 3 C/min from room temperature to 1160C. After the temperature reaches 1160C, it was held for 6 h to ensure a complete and homogeneous melting. The sample was then lowered with a rate of 1 cm/day to start the crystallization. The process continued for 4 days before the temperature was decreased down to room temperature with a rate of 3 C/min.
[…]
II.2 Magnetometry and neutron spectroscopy
Magnetization of Cu5SbO6 in an applied magnetic field oriented parallel ($\boldsymbol{H}_{ab}$) and perpendicular ($\boldsymbol{H}_{c^{*}}$) to the honeycomb plane was measured using a Quantum Design MPMS 3 vibrating-sample magnetometer. The temperature dependence was measured from $T=2$ to 400 K with $H=100$ Oe whereas the field dependence was measured from $H=0$ to $7\times 10^{4}$ Oe at $T=2$ and 120 K. For both field directions, an aligned crystal sample (see inset of Fig. 2(a)) with mass 9.48 mg was attached to a quartz rod whose magnetization was measured separately under the same conditions.
[…]
III Results and Discussion
III.1 Crystal structure and stacking faults
Figure 2: (a) A plot of $F_{\mathrm{obs}}$ against $F_{\mathrm{cal}}$ obtained from refinement of the single-crystal XRD data. The straight line is where $F_{\mathrm{obs}}=F_{\mathrm{cal}}$. The insets show a typical single-crystal sample and the $\theta$-$2\theta$ scan of a cleaved facet. (b) Powder XRD pattern of ground crystals. Open circles are observed intensity. The black line represents the calculated intensity based on the refined parameters. The blue line shows the difference between observed and calculated intensities. The green bars mark the position of Bragg peaks. (c) Normalized absorption spectra of CuO (orange) and Cu2O (green). The absorption coefficient $\mu x$ is shown on the left axis and the first derivative on the right axis. The asterisk and diamond symbols indicate the edge energy of Cu1+ and Cu2+ ions, respectively. (d) Normalized absorption spectrum of Cu5SbO6. The dashed line represents background of the spectrum. (e) Modulus of the Fourier transform of $\varphi(k)$ with the real part shown in (f). The open circles represent the data, and the red curve is the fitting with the model described in text. The dotted line is the fitting range (window).
Figure 3: (a) A projection of the crystal structure onto the $bc^{*}$ plane showing the ideal stacking sequence of Cu5SbO6. (b) Two equivalent representatives of in-plane translations of a honeycomb layer that contain one stacking fault depicted by the purple arrows. The grey arrows represent the ideal translation vector. ‘1’, ‘2’, and ‘3’ are sub-lattice labels of the honeycomb layer. The atom representations are the same as in Fig. 1(a). (c) A TEM image in the $ac$ plane of the sample synthesized at 1000 C. (d), (e) SAED patterns in the $(H,0,L)$ planes of samples synthesized at 1000 C and 1150 C, respectively. (f) A simulated SAED pattern computed with $p_{1}=p_{2}=1/2$. The crystal structure illustrations are generated by VESTA [23].
Table 1: Optimized atomic parameters of Cu5SbO6 obtained from refinements of the single-crystal and powder XRD data taken at room temperature. The unit cell is monoclinic with the space group $C2/c$ (no. 15). The lattice parameters are $a$ = 8.9291(2) Å, $b$ = 5.5974(1) Å, $c$ = 11.8529(2) Å, $\beta$ = 103.605(1). Numbers in parentheses are one standard deviation.
Atom Site $x/a$ $y/b$ $z/c$
Single-crystal x-ray diffraction
Sb $4c$ 3/4 3/4 0
Cu(1) $4e$ 0 0.6243(3) 1/4
Cu(2) $8f$ 0.8285(1) 0.0968(3) 0.2437(1)
Cu(3) $8f$ 0.4168(1) 0.7487(2) 0.00175(9)
O(1) $8f$ 0.4451(8) 0.118(1) 0.0912(6)
O(2) $8f$ 0.2227(8) 0.924(1) 0.9178(6)
O(3) $8f$ 0.6176(8) 0.381(1) 0.5969(6)
[…]
Powder x-ray diffraction
Sb $4c$ 3/4 3/4 0
Cu(1) $4e$ 0 0.6251(6) 1/4
Cu(2) $8f$ 0.8285(4) 0.0969(4) 0.2427(3)
Cu(3) $8f$ 0.4164(3) 0.749(1) 0.0019(3)
O(1) $8f$ 0.450(1) 0.112(2) 0.092(1)
O(2) $8f$ 0.225(1) 0.920(2) 0.916(1)
O(3) $8f$ 0.615(1) 0.377(2) 0.595(1)
$R_{\mathrm{p}}=6.42\%$, $R_{\mathrm{wp}}=9.05\%$, GoF = 1.87
Refinements of the crystal structure using the collected single-crystal and powder XRD data are based on the structural model reported in Ref [48]. The refinement with the single-crystal XRD data is performed with 799 unique reflections with $I_{\mathrm{obs}}>3\sigma$, where $I_{\mathrm{obs}}$ and $\sigma$ are the observed intensity and corresponding standard deviation, respectively, using jana2020 software package [53]. The refinement result is presented in Fig. 2(a) as a plot of observed structure factor $F_{\mathrm{obs}}$ against calculated structure factor $F_{\mathrm{cal}}$ computed from the refined atomic parameters listed in Table 1. As can be seen in Fig. 2(a), the model provides the values of the calculated structure factor that are fairly consistent with the observed data. The cleavage of a typical crystal was examined with the $\theta$-$2\theta$ scan and was found to be $(00L)$ planes as shown in the inset of Fig. 2(a). Fig. 2(b) displays the powder XRD pattern obtained from ground crystals. Comparing the position of the observed Bragg peaks with that derived from the model (green bars) indicates that impurities in the sample, if present, are negligible. The diffraction pattern is fitted to the model using the Rietveld method [54] as implemented in fullprof software suite [55]. The refinement yields the optimized lattice and atomic parameters listed in Table 1. The calculated diffraction pattern generated from the so-obtained parameters is illustrated in Fig. 2(b) as a black solid line which fits the observed data fairly well.
While XRD is one of the most powerful techniques for investigating the periodic structure of compounds, it is, however, practically incapable of distinguishing ions that have similar atomic factors as in the case of Cu1+ and Cu2+ in Cu5SbO6. Namely, while coordinates of Cu(1), Cu(2) (between honeycomb planes), and Cu(3) (in honeycomb planes) in the unit cell can be accurately identified, their oxidation states must be determined by another method. As Cu1+ and Cu2+ are quite similar in size, it is possible to have site interchange between them as in the case of Li3Cu2SbO6 in which considerable amounts of Li1+ and Cu2+ ions exchange sites [29, 30]. To confirm that all Cu(3) in Cu5SbO6 has 2+ oxidation state and, therefore, form full Cu2+ honeycomb layers, we turn to XAFS technique which is more effective for determining the oxidation state of a specific ion. The absorption coefficient, $\mu x$, of a material can be obtained from $I=I_{0}e^{-\mu x}\Leftrightarrow\mu x=-\ln(I/I_{0})$ where $I$ and $I_{0}$ are transmitted and incident intensities, respectively. Fig. 2(c) and 2(d) show the collected XAFS spectra of CuO, Cu2O and Cu5SbO6, respectively. Inspection on the results in Fig. 2(c) shows that Cu1+ ions in Cu2O (green) and Cu2+ ions in CuO (orange) have unequal edge energies. Estimated values of edge $E_{0}$ for Cu1+ and Cu2+ ions are determined from their absorption spectra as the energy $E$ at which the first derivative of $\mu x$ is maximized, which is done using the athena module in the demeter software suite [56]. The first derivatives of $\mu x(E)$ for both Cu2O and CuO are displayed as dashed lines in Fig. 2(c) with the asterisk and diamond symbols marking the estimated $E_{0}$ of Cu1+ and Cu2+ ions, respectively. The difference in $E_{0}$ of the two ions is about 10 eV.
To determine the oxidation states of Cu ions in Cu5SbO6, we performed the extended x-ray absorption fine structure (EXAFS) analysis. EXAFS analysis is performed in the oscillating region of the spectrum [57] which is quantified by $\varphi(k)\equiv\mu x-\mu_{0}x$ where $\mu_{0}x$ is the background signal (dashed line in Fig. 2(d)) and $k$ is related to $E$ through $k=\sqrt{2m_{\mathrm{e}}(E-E_{0})}/\hbar$. Here $m_{\mathrm{e}}$ is the electron mass. Contrary to XRD that depends upon average periodic structures, EXAFS relies on local environments around a particular ion. As the local environment of Cu sites within the honeycomb planes and that in between the planes are completely different, each site, therefore, produces a distinct oscillating pattern in $\varphi(k)$. To construct a model for the fitting, the crystal structure listed in Table 1 is used to generate all possible electron scattering paths around each Cu site, and $\varphi(k)$ of Cu5SbO6 is then a superposition of $\varphi_{i}(k)$’s from paths having significant amplitudes. To facilitate the fitting, the observed spectrum is Fourier-transformed to a real-space variable $R$, whose modulus and real part are illustrated in Figs. 2(e) and 2(f), respectively. The fitting is performed using the artemis module in the in demeter software suite [56]. The fitting parameters include path distance corrections $\delta R$ along with thermal deviations $\sigma$, Cu amplitude reduction factor $S_{0}^{2}$, and edge corrections $\delta E_{0}$. The amplitude reduction factor $S_{0}^{2}$ for Cu atom is obtained separately from fitting $\varphi(R)$ collected with a Cu foil sample and is found to be 0.89(5). The initial values of $E_{0}$ for each Cu site are set to the value of Cu1+ ion, namely, the value marked by the asterisk in Fig. 2(c). The fitting result is shown as red curves in Figs. 2(e) and 2(f) with $R=0.262\%$. The optimized values of $\delta E_{0}$ for Cu(1) and Cu(2) are found to be the same (and, therefore, are constrained to the same value in the final refinement) which equals to 3.0(5) eV, while $\delta E_{0}$ of Cu(3) is 13.2(3) eV. Hence, $E_{0}$ of Cu(3) residing in the honeycomb planes is larger than $E_{0}$ of the other two sites in between the planes by about 10 eV, i.e., the oxidation state of Cu(3) is 2+ and that of Cu(1) and Cu(2) is 1+.
[…]
A selected-area electron diffraction (SAED) pattern for the powder sample synthesized at 1000 C is shown in Fig. 3(d) in which diffusive scattering streaks are observed along $(2,0,L)$ and $(4,0,L)$ but are absent along $(0,0,L)$ and $(6,0,L)$. The streaks observed reflect lack of periodicity in the $c^{*}$ direction pointing to translations of the Cu2+ honeycomb layers that intermittently intervene the ideal ‘123’ sequence. To gain more insight into this mis-translation, we next investigate a TEM image of the $ac$ plane of this sample as displayed in Fig. 3(c). The image clearly features an irregularity in the stacking pattern and hints at the presence of two in-plane translations along the $a$ direction. Inspecting Fig. 3(b) suggests that, in addition to the ideal in-plane translation, there is another possible translation (depicted by purple arrows) of the Cu2+ honeycomb layers that translates Sb5+ from sub-lattice ‘1’ to ‘3’, ‘3’ to ‘2’, and ‘2’ to ‘1’. Thus, an appropriate model for stacking faults in Cu5SbO6 corresponds to intervening the ideal ‘123’ sequence by a stacking-fault translation at certain frequency. It is worth noting that the ideal in-plane translation vector can be represented by (1/3, 0, 0) and that of the stacking fault by $(2/3,0,0)$ or, equivalently, (1/3,0,0)(-1/3,0,0). This model immediately provides a description for the disappearance of diffusive streaks along $(6,0,L)$. As the stacking-fault translation vector is $(2/3,0,0)$ which can be obtained by adding $(1/3,0,0)$ to the ideal translation vector. The additional phase factor caused by the extra $(1/3,0,0)$ results in a phase difference of $2\pi$ for $(6,0,L)$ planes, hence, constructive interference. The streaks are not observed along $(0,0,L)$ since the phase factor due to these planes always adds in-phase irrespective of the in-plane translations. To further verify the model, a SAED pattern is simulated using faults software package [60, 61, 62]. The frequency of the ideal $p_{1}$ and of the stacking-fault translations $p_{2}$ are introduced which are subjected to a constraint $p_{1}+p_{2}=1$. Fig. 3(f) illustrates a simulated SAED pattern of an extreme case where $p_{1}=p_{2}=1/2$ which can well reproduce the diffusive intensities along $(2,0,L)$ and $(4,0,L)$ in the observed SAED pattern in Fig. 3(d). In addition, the model can also reproduce the broad peak in the XRD data of a powder sample as illustrated in the inset of Fig. 5(a) of Ref. [48].
A SAED pattern for the sample synthesized at 1150 C, a temperature just below the melting point, shows a strikingly different result as shown in Fig. 3(e). In this case, the diffusive scattering streaks are unobserved reflecting a restoration of the periodicity along the $c^{*}$ direction. The absence of the diffusive streaks in this SAED pattern is consistent with the absence of peak broadening, especially in $19^{\circ}<2\theta<23^{\circ}$, in the powder XRD data collected from ground crystals (see Fig. 2(b)). The improvement of stacking faults in Cu5SbO6 by raising the synthesis temperature was also reported in Ref. [48].
III.2 Magnetization
Figure 4: (a) Magnetic susceptibility as a function of temperature, $\chi(T)$, measured with an applied magnetic field $H=100$ Oe. The solid curves are the best fits. (b), (c) Magnetization of dimerized spins $M_{\mathrm{dim}}$ as a function of applied magnetic field normalized with the saturated magnetization $M_{\mathrm{sat}}=g\mu_{\mathrm{B}}S$ measured at $T=120$ K. The solid lines are model calculations described in text. The data are collected with $\boldsymbol{H}$ parallel to the $ab$ plane (circle symbols) and to the $c^{*}$ axis (square symbols).
The temperature dependence of magnetic susceptibility, $\chi=M/H$, for the in-plane ($\boldsymbol{H}_{ab}$) and out-of-plane ($\boldsymbol{H}_{c^{*}}$) field directions are shown in Fig. 4(a). The holder’s susceptibility measured separately is found to be smaller than the measurement uncertainty of the sample’s susceptibility and is, therefore, neglected. As the temperature decreases from 400 K, a broad peak indicative of short-range correlations develops at 120 K followed by an upturn below 20 K. No signature of thermal phase transitions is observed down to 2 K. Regardless of the upturn, the absence of anomalies along with the gradual decrease in $\chi(T)$ toward zero suggests the formation of spin dimers in Cu5SbO6 at low temperatures. To further investigate the upturn below 20 K, magnetization curves, $M(H)$, for both field directions were measured at 2 K where the signal from dimerized spins can be safely ignored. It is found that the $M(H)$ curves exhibit a Brillouin function that resembles Curie paramagnetism and, therefore, points to the presence of quasi-free spins in the sample under study. Hence, an appropriate model of the susceptibility can be expressed as
$\chi(T)=f_{\mathrm{free}}\chi_{\mathrm{free}}(T)+f_{\mathrm{dim}}\chi_{\mathrm{dim}}(T)+\chi_{0}$ (1)
where $f_{\mathrm{free}}$ and $f_{\mathrm{dim}}$ are population fractions of quasi-free and dimerized spins, respectively, with a constraint $f_{\mathrm{free}}+f_{\mathrm{dim}}=1$. $\chi_{0}$ is a temperature-independent background susceptibility from other non-magnetic ions in Cu5SbO6 and core shells of Cu2+. $\chi_{\mathrm{free}}(T)=C/T$ with the Curie constant expressed as $C=N_{\mathrm{A}}p^{2}\mu^{2}_{\mathrm{B}}/(3k_{\mathrm{B}})$ where $N_{\mathrm{A}}$ is Avogadro constant, $\mu_{\mathrm{B}}$ is Bohr magneton, $k_{\mathrm{B}}$ is Boltzmann constant, and $p=g\sqrt{S(S+1)}$ is the effective Bohr magneton number with $g$ being Landé factor of quasi-free spins.
We determine $C$ and $f_{\mathrm{free}}$ from the $M(H)$ curves for both field directions measured at $T=2$ K. To model the $M(H)$ curve at this temperature, two terms responsible for paramagnetic and background magnetization are required. These two terms are $M_{\mathrm{free}}(H)\propto B_{1/2}(H)$ where $B_{1/2}(H)$ is a Brillouin function for $S=\frac{1}{2}$, and $M_{0}(H)\propto H$. Therefore,
$M(H)=\frac{1}{2}g\mu_{\mathrm{B}}N_{\mathrm{free}}\tanh\left(\frac{g\mu_{\mathrm{B}}H}{2k_{\mathrm{B}}T}\right)+M_{0}(H),$ (2)
[…]
We perform quantum Monte Carlo (QMC) simulations to acquire a model for $\chi_{\mathrm{dim}}(T)$. The reduced susceptibility, $\chi^{*}(t)$, for $0.01\leq t\leq 10$ where $t=k_{\mathrm{B}}T/J_{\mathrm{max}}$ is calculated using the loop algorithm [63, 64] implemented in the alps simulation package [65]. The simulation is performed for the $J_{1}$-$J_{2}$-$J_{4}$ model with $J_{\mathrm{max}}=J_{1}=16.5$ meV, $J_{2}=-6$ meV, and $J_{4}=3$ meV. The determination of these coupling parameters will be discussed in the next section. The simulations are performed on a cluster of $\times100$ spins with a periodic boundary condition using 100 000 Monte Carlo steps for thermalization and 500 000 Monte Carlo steps after thermalization. To obtain a functional form of $\chi^{*}(t)$, we fit simulation results with [66]
$\chi^{*}(t)=\frac{\exp(-\Delta^{*}/t)}{4t}\mathcal{P}^{(q)}_{(r)}(t),$ (3)
where $\mathcal{P}^{(q)}_{(r)}(t)$ is the Padé approximant expressed as
$\mathcal{P}^{(q)}_{(r)}(t)=\frac{1+\sum_{n=1}^{q}N_{n}/t^{n}}{1+\sum_{n=1}^{r}D_{n}/t^{n}}.$
Here $\Delta^{*}$, $N_{n}$, and $D_{n}$ are fitting parameters. We have found that the Padé approximant with $q=r=5$, i.e., $\mathcal{P}^{(5)}_{(5)}(t)$, yields a sufficient accuracy with the deviation between the simulation results and equation (3) in the order of $10^{-3}$. $\chi_{\mathrm{dim}}(T)$ is related to $\chi^{*}(t)$ by
$\chi_{\mathrm{dim}}(T)=\frac{N_{\mathrm{A}}g^{2}\mu_{\mathrm{B}}^{2}}{J_{\mathrm{max}}}\chi^{*}\left(\frac{k_{\mathrm{B}}T}{J_{\mathrm{max}}}\right).$ (4)
[…]
To reaffirm the optimized parameters, we further perform quantum Monte Carlo (QMC) simulations to calculate the reduced magnetization, $M^{*}(h)$, with the same set of coupling parameters for $0\leq h\leq 7\times 10^{-2}$ where $h=g\mu_{\mathrm{B}}H/J_{\mathrm{max}}$ and $T=120$ K using the directed-loop [67, 68] algorithm. The calculation conditions are the same as for $\chi^{*}(t)$. $M_{\mathrm{dim}}(H)$ is related to $M^{*}(h)$ by
$M_{\mathrm{dim}}(H)={}$ $g\mu_{\mathrm{B}}M^{*}\left(\frac{g\mu_{\mathrm{B}}H}{J_{\mathrm{max}}}\right).$
The simulated $M_{\mathrm{dim}}(H)$ is calculated with the optimized Landé factor $g$ and then compared with the background-corrected experimental data. As seen in Figs. 4(b) and 4(c), a reasonable agreement is achieved between the data and model calculations confirming the validity of the Landé factors and the exchange couplings.
III.3 Triplet excitations
Figure 5: (a) The observed powder-averaged magnetic intensity map at $T=4$ K (b) The calculated powder-averaged magnetic intensity map using the $J_{1}$-$J_{2}$-$J_{4}$ model with $J_{1}=16.5$ meV, $J_{2}=-6$ meV, and $J_{4}=3$ meV. (c) The density of states (DOS) of the $J_{1}$-$J_{2}$-$J_{3}$ model and of (d) the $J_{1}$-$J_{2}$-$J_{4}$ model.
Figure 6: (a) $E$-scan data with $Q=1.97\;\mathrm{\AA }^{-1}$ at 2.8 K. Solid lines are best fits to the model described in text. (b) Temperature dependence of $\hbar\omega$ for $M_{1}$ (solid squares) and $M_{2}$ (open squares) peaks. Dashed lines are guides to the eyes. Error bars represent one standard deviation.
Data reduction and powder averaging of the TOF INS data collected from the AMATERAS spectrometer were performed using the utsusemi software package [52]. The obtained intensity maps at $T=4$ K and 200 K are displayed in Figs. S.1(a) and S.1(b), respectively, in the Supplemental Information (SI). Also included in Fig. S.1(a) is the excitation energy (circle symbols) extracted from the $E$-scan data taken from the BT-7 spectrometer at $T=2.8$ K. In Fig. S.1(b), at $T=200$ K where the magnetic correlations are expected to subside (see also Fig. 6(a)), there remain excitations whose intensity noticeably increases with the magnitude of the wavevector $Q$ and temperature. We, therefore, attribute these excitations to phonons. To obtain the magnetic scattering intensity map, $I_{\mathrm{mag}}(Q,\omega)$, as shown in Fig. 5(a), we subtracted the data at the two temperatures after rescaling the data at 200 K by the temperature-dependent factor $1+n(T)=(1-\exp(-E/k_{\mathrm{B}}T)^{-1}$ where $n(T)$ is the Bose factor. It should be noted that the scattering intensity seen at $\hbar\omega=12$ meV and $Q>3$ Å-1 is most likely due to residual phonon intensity. To acquire the key characteristics of the magnetic scattering, $I_{\mathrm{mag}}(Q,\omega)$ is then integrated over the wavevector $1\;\mathrm{\AA ^{-1}}\leq Q\leq 3\;\mathrm{\AA ^{-1}}$. The integrated intensity shown in Fig. S.2(b) features two peaks at $\hbar\omega=15$ meV and 18 meV and covers an extensive range from the band bottom at $\sim$ 11 meV to the band top at $\sim$ 21 meV amounting to a sizable bandwidth of $\sim$ 10 meV. No magnetic excitation is detected at $\hbar\omega\lesssim 11$ meV or $\hbar\omega\gtrsim 21$ meV.
To explore the temperature evolution of the excitation spectrum, we next examine the TA $E$-scan data collected at base and elevated temperatures from the BT-7 spectrometer. Fig. 6(a) shows the data measured at $2.8\;\mathrm{K}\leq T\leq 200\;\mathrm{K}$ with a fixed wavevector $Q=1.96\;\mathrm{\AA }^{-1}$. At $T=2.8$ K, the data exhibit three excitation peaks at 12 meV, 15 meV, and 18 meV, denoted by $P$, $M_{1}$, and $M_{2}$, respectively. At elevated temperatures, the $P$ peak becomes stronger whereas the $M_{1}$ and $M_{2}$ peaks weaken and eventually disappear at 200 K. Therefore, the $M_{1}$ and $M_{2}$ peaks are consistent with the two peaks observed in the integrated magnetic intensity shown in Fig. S.2(b). For the $P$ peak, however, it is essential to note that in the TOF data (see Figs. S.1(a) and S.1(b) in the SI) there exist a phonon band as well as magnetic scattering at $\hbar\omega=12$ meV. Therefore, the $P$ peak, especially at $T=2.8$ K, results from the superposition of these two sources. As the temperature is elevated, the magnetic contribution diminishes leaving only the enhanced phonon contribution. We fit the spectra to a Lorentzian function combined with a linear background to acquire the excitation energy, $\hbar\omega$, for each $T$ and $Q$. The width of the $M_{1}$ and $M_{2}$ peaks are assumed identical. Fitting the data (i) at $T=2.8$ K and with $1.12\;\mathrm{\AA }^{-1}\leq Q\leq 3.09\;\mathrm{\AA }^{-1}$ and (ii) at $2.8\;\mathrm{K}\leq T\leq 200\;\mathrm{K}$ with $Q=1.96\;\mathrm{\AA }^{-1}$ yields $\hbar\omega$ plotted as circle symbols in Fig. S.1(a) and a $T$-dependent $\hbar\omega$ shown in Fig. 6(b). In alignment with $\chi(T)$ data, the monotonic increase in $\hbar\omega(T)$ points out that the gapped excitation observed in Cu5SbO6 is a triplet excitation (or triplon) since, in a dimerized magnet, triplons are subjected to the hard-core constraint that prevents them from occupying the same dimer site resulting in repulsion between them [69, 70, 71, 72] as opposed to long-range ordered magnets for which the gap decreases with increasing temperature due to the reduction in ordered moment.
To acquire an appropriate spin Hamiltonian for Cu5SbO6, we employ the first-order dimer expansion to derive the energy dispersion, $\Delta E_{\boldsymbol{Q}}$, of triplet excitations for both $J_{1}$-$J_{2}$-$J_{3}$ and $J_{1}$-$J_{2}$-$J_{4}$ models (a detailed derivation can be found in the Appendix). The expressions for the dispersion of both models are given in Eqs. (14) and (15). In these expressions, the wavevector $\boldsymbol{Q}$ is written in terms of the basis vectors $\boldsymbol{b}_{1}$, $\boldsymbol{b}_{2}$, and $\boldsymbol{b}_{3}$ defined in Eqs. (8)-(10). Here, $q_{i}$ represents the component of each basis vector. Let us begin with the $J_{1}$-$J_{2}$-$J_{3}$ model with the dispersion given in Eq. (14). For an isolated FM-AFM chain, the dispersion along the $q_{1}$ direction oscillates around the energy set by the intra-dimer AFM $J_{1}$ coupling with the amplitude determined by the magnitude of the inter-dimer FM $J_{2}$ coupling. The incorporation of the interchain AFM $J_{3}$ coupling predominantly only results in a modulation of the top of the dispersion along the $q_{2}$ direction (see Figs. 8(a)-8(d)). To see how the dispersion produces the $M_{1}$ and $M_{2}$ peaks in the magnetic scattering, we consider the density of states (DOS), $D(\omega)$, as defined by
$D(\omega)\propto\int d\boldsymbol{Q}\;\delta(\hbar\omega-\Delta E_{\boldsymbol{Q}}),$
where the integration is over a unit cell of the reciprocal lattice formed by dimers. Approximating the delta function by a Lorentzian function, $\delta(x)=\epsilon/\pi(x^{2}+\epsilon^{2})$, where $\epsilon$ is a small positive number, we finally obtain DOS for the $J_{1}$-$J_{2}$-$J_{3}$ model as shown in Fig. 5(c). The result is presented for four different values of the interchain coupling $J_{3}$ with $J_{1}=16.95$ meV, $J_{2}=-3.9$ meV. When $J_{3}=0$, two peaks appear at the top and bottom of the DOS. As $J_{3}$ increases, the peak at the top shifts towards lower energy and the tail develops towards higher energy. Although $J_{3}\approx 1$ meV correctly reproduces peaks in DOS at 15 meV and 18 meV, this model does not provide a tail below 15 meV regardless of the value of $J_{3}$ which is in contradiction to the observed spectrum. Therefore, the $J_{1}$-$J_{2}$-$J_{3}$ model is not viable in explaining the data.
We next turn to the $J_{1}$-$J_{2}$-$J_{4}$ model. The introduction the interlayer AFM $J_{4}$ coupling to the system of isolated FM-AFM chains delivers a slightly distinct effect on the dispersion. Unlike with $J_{3}$, it is found that modulation along the $q_{3}$ direction appears at both the top and the bottom of the dispersion (see Figs. 8(e)-8(h)). To obtain DOS for this model, we first consider the dispersion in Eq. (15) from which the energy of $M_{1}$ and $M_{2}$ peaks, denoted by $E_{1}$ and $E_{2}$, respectively, are given by
$E_{1}={}$ $J_{1}-\frac{|J_{2}|}{2}+\frac{|J_{4}|}{2},$
$E_{2}={}$ $J_{1}+\frac{|J_{2}|}{2}-\frac{|J_{4}|}{2},$
which give $2J_{1}=E_{1}+E_{2}$ and $|J_{2}|=|J_{4}|+(E_{2}-E_{1})$. Using $E_{1}=15.0(1)$ meV and $E_{2}=18.0(2)$ meV, we can identify $J_{1}=16.5(1)$ meV and $|J_{2}|=|J_{4}|+3.0$ meV. Fig. 5(d) shows DOS of the $J_{1}$-$J_{2}$-$J_{4}$ model calculated with four different values of $|J_{4}|$. With these coupling parameters the DOS correctly reproduces peaks at 15 meV and 18 meV. Moreover, in contrast to the previous model, as $J_{4}$ increases, tails in DOS develop at both higher and lower energy which resembles the pattern observed in the INS spectrum. Based on the INS data, one can obtain an estimate of $|J_{4}|\approx 3$ meV, and, thus, $|J_{2}|\approx 6$ meV. It is essential to note that in the presence of stacking faults, although the magnitude of $J_{4}$ might be unchanged as the path responsible for the coupling still involves the same ions and geometry, the dimer connectivity is different and the $J_{1}$-$J_{2}$-$J_{4}$ model might be inapplicable.
It is natural to extend the model to include the $J_{3}$ coupling in the $J_{1}$-$J_{2}$-$J_{4}$ model. To this end, let us examine the effect of $J_{3}$ on the current model by calculating the DOS of the extended model with different values of $J_{3}$ as described in the SI. As can be seen in Fig. S.2(c), the presence of the $J_{3}$ coupling results in a suppression of the 18 meV peak and a further elongation of the tail above 21 meV. These effects become more enhanced as $J_{3}$ grows. Comparison with the integrated magnetic intensity in Fig. S.2(b) suggests an upper bound for $J_{3}$ of merely $\sim 0.3$ meV, which can safely be neglected when taking the uncertainty of the coupling parameters into account. Finally, to verify the $J_{1}$-$J_{2}$-$J_{4}$ model, a powder-averaged INS spectrum was simulated using the proposed set of coupling parameters as outlined in the SI. The simulated result shown in Fig. 5(b) compares very well with the observed spectrum justifying the $J_{1}$-$J_{2}$-$J_{4}$ model as adequate and appropriate for Cu5SbO6.
Future INS studies on high-quality single-crystal samples may shed more light on some exotic aspects of the triplon excitations and allow the spin-spin correlations to be mapped out more precisely. The latter, as discussed in Refs. [73, 74, 75], play a crucial role in determining the degree of hidden quantum entanglement in a many-body system such as a spin chain. It has been argued in these studies that the spatial extent of quantum entanglement is directly related to the spin-spin correlations that can be measured experimentally by means of the neutron scattering technique. In this arena, as pointed out in Ref. [22], an alternating FM-AFM Heisenberg spin chain may triumph over its alternating AFM-AFM counterpart owing to the non-zero spin-spin correlations on both FM and AFM bonds. Given this, Cu5SbO6 can serve as a model system for exploring the spatial evolution of quantum entanglement manifesting in a many-body system.
IV Conclusion
We employed x-ray spectroscopy and electron microscopy to refine the crystal structure and the stacking fault model of Cu5SbO6. Based on the crystal structure, Cu5SbO6 can accommodate alternating FM-AFM dimerized spin chains similar to the related compounds with the FM bond forming between Cu2+ ions within the Cu2O6 double plaquettes and the AFM bond, hence, the dimers, forming across the hexagon in the honeycomb planes. However, in contrast to the related compounds, the inter-chain couplings that arise between the honeycomb planes can be significant for Cu5SbO6 owing to the different geometry of the super-exchange paths that enhance orbital overlap. Stacking faults present in powder samples are a result of two inequivalent in-plane translations, i.e., (1/3, 0, 0) and (2/3, 0, 0), that interchangeably shift the honeycomb planes along the $a$-direction with a frequency depending on the synthesis temperature.
Based on a first-order dimer expansion approximation, we showed that, unlike the related compounds, the $J_{1}$-$J_{2}$-$J_{3}$ model cannot be a complete description for the magnetic properties of Cu5SbO6. With a specific set of coupling parameters, the density of states of such a model, whereas containing peaks at 15 meV and 18 meV, does not exhibit a tail below 15 meV in the INS spectrum. On the other hand, the $J_{1}$-$J_{2}$-$J_{4}$ model is capable of reproducing the entire spectrum and, therefore, proves to be a more appropriate starting point for the spin Hamiltonian for Cu5SbO6. The coupling parameters of the $J_{1}$-$J_{2}$-$J_{4}$ model estimated from the powder-averaged INS data are $J_{1}=16.5(1)$ meV, $|J_{2}|\approx 6\;\mathrm{meV}$, and $|J_{4}|\approx 3\;\mathrm{meV}$. With the obtained coupling parameters, we further tested the model by simulating a temperature-dependent susceptibility and field-dependent magnetization. Fitting the simulated susceptibility to the experimental data returns sensible Landé factors, i.e., $g_{ab}=2.177(2)$ and $g_{c^{*}}=2.309(2)$, for Cu2+ moments in plaquette environments. Moreover, the simulated magnetization curves calculated with the optimized Landé factors compare reasonably well with the experimental data. Considering these results, Cu5SbO6 may be included in a rare collection of compounds that host interacting FM-AFM quantum spin chains which can serve as a playground for exploring novel emergent quantum collective phenomena.
Acknowledgements.
This research has received funding support from the NSRF via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation (PMU-B) (Grant No. B13F680079). Work at Mahidol University was partly supported by the Thailand Center of Excellence in Physics and Mahidol University (Strategic Research Fund): fiscal year 2024. C.P. is grateful for the support by the DPST scholarship from the Institute for the Promotion of Teaching Science and Technology. C.P. and K. Matan thank Dr. Khetpakorn Chakarawet from the Department of Chemistry, Mahidol University, for his assistance in single-crystal XRD measurements. C.P. and K. Matan thank Masaki Ageishi for his assistance in TEM experiments at Tohoku University, Japan. W.-T.C. acknowledges the support of the National Science and Technology Council in Taiwan (Grants No. 113-2124-M-002-006 and 114-2124-M-002-010) and the Featured Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education in Taiwan (Grants No. 113L9008). K. Morita was supported by JSPS KAKENHI (Grant No. JP25K07230). The experiment on AMATERAS was performed with the approval of J-PARC (Proposal No. 2019A0194 (R)). We are grateful for the support from the National Institute of Standards and Technology, U.S. Department of Commerce, for providing neutron research facilities used in this work. The identification of any commercial product or trade name does not imply endorsement or recommendation by the National Institute of Standards and Technology. Synchrotron Light Research Institute (Public Organization), Thailand, is acknowledged for the provision of XAS beamtime.
Appendix A Dimer expansion for triplet excitations
Figure 7: A lattice formed by dimers depicted as green ovals for the $J_{1}$-$J_{2}$-$J_{3}$ and $J_{1}$-$J_{2}$-$J_{4}$ models are shown in (a) and (b), respectively. $i$ is the dimer identifier in the lattice, and $\nu=0,1$ is the spin identifier in the dimer. $\boldsymbol{a}_{1}$, $\boldsymbol{a}_{2}$, and $\boldsymbol{a}_{3}$ are the primitive translation vectors.
In this appendix, the dispersion relation as well as the dynamical structure factor for the $J_{1}$-$J_{2}$-$J_{3}$ and $J_{1}$-$J_{2}$-$J_{4}$ models are derived. Fig. 7 shows the lattices formed by dimers for the two models where the primitive translation vectors $\boldsymbol{a}_{1}$, $\boldsymbol{a}_{2}$, $\boldsymbol{a}_{3}$ are expressed as
$\boldsymbol{a}_{1}={}$ $l_{1}\boldsymbol{e}_{x},$ (5)
$\boldsymbol{a}_{2}={}$ $l_{2}(\cos\phi_{\mathrm{cso}}\boldsymbol{e}_{x}+\sin\phi_{\mathrm{cso}}\boldsymbol{e}_{y}),$ (6)
$\boldsymbol{a}_{3}={}$ $l_{3}(\cos\theta_{\mathrm{cso}}\boldsymbol{e}_{x}+\sin\theta_{\mathrm{cso}}\boldsymbol{e}_{z}).$ (7)
$\boldsymbol{a}_{1}$ and $\boldsymbol{a}_{2}$ form the dimer lattice of the $J_{1}$-$J_{2}$-$J_{3}$ model whereas $\boldsymbol{a}_{1}$ and $\boldsymbol{a}_{3}$ form that of the $J_{1}$-$J_{2}$-$J_{4}$ model. The primitive translation vectors of the reciprocal dimer lattice $\boldsymbol{b}_{1}$, $\boldsymbol{b}_{2}$, $\boldsymbol{b}_{3}$ are
$\boldsymbol{b}_{1}={}$ $\frac{2\pi}{l_{1}}(\boldsymbol{e}_{x}-\cot\phi_{\mathrm{cso}}\boldsymbol{e}_{y}-\cot\theta_{\mathrm{cso}}\boldsymbol{e}_{z}),$ (8)
$\boldsymbol{b}_{2}={}$ $\frac{2\pi}{l_{2}}\csc\phi_{\mathrm{cso}}\boldsymbol{e}_{y},$ (9)
$\boldsymbol{b}_{3}={}$ $\frac{2\pi}{l_{3}}\csc\theta_{\mathrm{cso}}\boldsymbol{e}_{z}.$ (10)
$\boldsymbol{b}_{1}$ and $\boldsymbol{b}_{2}$ form the reciprocal dimer lattice of the $J_{1}$-$J_{2}$-$J_{3}$ model whereas $\boldsymbol{b}_{1}$ and $\boldsymbol{b}_{3}$ form that of the $J_{1}$-$J_{2}$-$J_{4}$ model. Note that for Cu5SbO6, $l_{1}=9.00$ Å, $l_{2}=5.30$ Å, $l_{3}=6.55$ Å, $\theta_{\mathrm{cso}}=62^{\circ}$, $\phi_{\mathrm{cso}}=30^{\circ}$.
A.1 Energy dispersion
In our Hamiltonian $H=H_{0}+H^{\prime}$ for the two models, the common unperturbed term $H_{0}$ is defined by
$H_{0}=J_{1}\sum_{i}\boldsymbol{S}_{i,0}\cdot\boldsymbol{S}_{i,1},$
where $i$ is an identifier given to a dimer and $\boldsymbol{S}_{i,\nu}$ is the $\nu$-th spin operator on the $i$-th dimer (see Fig. 7). The perturbation term $H^{\prime}$ for each model is given by
$H_{123}^{\prime}={}$ $J_{2}\sum_{i}\boldsymbol{S}_{i,1}\cdot\boldsymbol{S}_{i+\boldsymbol{a}_{1},0}$
$+{}$ $J_{3}\sum_{i}\boldsymbol{S}_{i,1}\cdot(\boldsymbol{S}_{i+\boldsymbol{a}_{2},0}+\boldsymbol{S}_{i+\boldsymbol{a}_{1}-\boldsymbol{a}_{2},0}),$
$={}$ $\sum_{i}\sum_{\rho=\boldsymbol{a}_{1},\boldsymbol{a}_{2},\boldsymbol{a}_{1}-\boldsymbol{a}_{2}}J_{\rho}\boldsymbol{S}_{i,1}\cdot\boldsymbol{S}_{i+\rho,0}$
$H_{124}^{\prime}={}$ $J_{2}\sum_{i}\boldsymbol{S}_{i,1}\cdot\boldsymbol{S}_{i+\boldsymbol{a}_{1},0}$
$+{}$ $J_{4}\sum_{i}\boldsymbol{S}_{i,1}\cdot\boldsymbol{S}_{i+\boldsymbol{a}_{1}-\boldsymbol{a}_{3},0},$
$={}$ $\sum_{i}\sum_{\rho=\boldsymbol{a}_{1},\boldsymbol{a}_{1}-\boldsymbol{a}_{3}}J_{\rho}\boldsymbol{S}_{i,1}\cdot\boldsymbol{S}_{i+\rho,0}.$
The wave function of the singlet sea $|s\rangle$ is expressed as
$|s\rangle=\prod_{n}|s\rangle_{n},\;\mathrm{where}\;|s\rangle_{n}\equiv\frac{1}{\sqrt{2}}(|\uparrow\downarrow\rangle_{n}-|\downarrow\uparrow\rangle_{n}).$
Here $|\uparrow\downarrow\rangle_{n}$, $|\downarrow\uparrow\rangle_{n}$, … represent the state of the $n$-th dimer in the Ising basis. We also define the single-triplet-dimer state $|t^{+}_{l}\rangle$ (STD), in which only the $l$-th dimer in the singlet sea is in $|t^{+}\rangle_{l}\equiv|\uparrow\uparrow\rangle_{l}$, as
$|t^{+}_{l}\rangle=|t^{+}\rangle_{l}\prod_{n(\neq l)}|s\rangle_{n}.$
Multiplying $H_{0}$ by $|t^{+}_{l}\rangle$, we immediately get
$H_{0}|t^{+}_{l}\rangle=J_{1}\sum_{i}(\boldsymbol{S}_{i,0}\cdot\boldsymbol{S}_{i,1})|t^{+}_{l}\rangle=J_{1}\left(-\frac{3}{4}N+1\right)|t^{+}_{l}\rangle$.
Next, we multiply $H^{\prime}$ by $|t^{+}_{l}\rangle$. Notice that each exchange operator in the perturbation term involves two dimers. In order to evaluate $H^{\prime}|t^{+}_{l}\rangle$, we explicitly calculate the following three patterns of terms for the two dimers. The first one is $\boldsymbol{S}_{l,1}\cdot\boldsymbol{S}_{l+\rho,0}|t^{+}\rangle_{l}|s\rangle_{l+\rho}$, which can be evaluated as
$\boldsymbol{S}_{l,1}\cdot\boldsymbol{S}_{l+\rho,0}|t^{+}\rangle_{l}|s\rangle_{l+\rho}={}$ $\left[S^{z}_{l,1}S^{z}_{l+\rho,0}+\frac{1}{2}\left(\underbrace{S^{+}_{l,1}}_{\to 0}S^{-}_{l+\rho,0}+S^{-}_{l,1}S^{+}_{l+\rho,0}\right)\right]|t^{+}\rangle_{l}|s\rangle_{l+\rho}$
$={}$ $\frac{1}{4}|t^{+}\rangle_{l}|t^{0}\rangle_{l+\rho}-\frac{1}{2}\frac{1}{\sqrt{2}}\left(|t^{0}\rangle_{l}+|s\rangle_{l}\right)\frac{1}{\sqrt{2}}|t^{+}\rangle_{l+\rho}$
$={}$ $-\frac{1}{4}|s\rangle_{l}|t^{+}\rangle_{l+\rho}+\frac{1}{4}|t^{+}\rangle_{l}|t^{0}\rangle_{l+\rho}-\frac{1}{4}|t^{0}\rangle_{l}|t^{+}\rangle_{l+\rho},$
which gives
$\boldsymbol{S}_{l,1}\cdot\boldsymbol{S}_{l+\rho,0}|t^{+}_{l}\rangle=-\frac{1}{4}|t^{+}_{l+\rho}\rangle+\textnormal{(non-STD states)}.$
The second one is $\boldsymbol{S}_{l-\rho,1}\cdot\boldsymbol{S}_{l,0}|s\rangle_{l-\rho}|t^{+}\rangle_{l}$, which can be evaluated as
$\boldsymbol{S}_{l-\rho,1}\cdot\boldsymbol{S}_{l,0}|s\rangle_{l-\rho}|t^{+}\rangle_{l}=$ $-\frac{1}{4}|t^{+}\rangle_{l-\rho}|s\rangle_{l}-\frac{1}{4}|t^{0}\rangle_{l-\rho}|t^{+}\rangle_{l}$
$+\frac{1}{4}|t^{+}\rangle_{l-\rho}|t^{0}\rangle_{l},$
which gives
$\boldsymbol{S}_{l-\rho,1}\cdot\boldsymbol{S}_{l,0}|t^{+}_{l}\rangle=-\frac{1}{4}|t^{+}_{l-\rho}\rangle+\textnormal{(non-STD states)}.$
The final one is $\boldsymbol{S}_{i,1}\cdot\boldsymbol{S}_{i+\rho,0}|s\rangle_{i}|s\rangle_{i+\rho}$ with $i\neq l$ and $i+\rho\neq l$, which can be evaluated as
$\boldsymbol{S}_{i,1}\cdot\boldsymbol{S}_{i+\rho,0}|s\rangle_{i}|s\rangle_{i+\rho}=$ $-\frac{1}{4}|t^{0}\rangle_{i}|t^{0}\rangle_{i+\rho}+\frac{1}{4}|t^{+}\rangle_{i}|t^{-}\rangle_{i+\rho}$
$-\frac{1}{4}|t^{-}\rangle_{i}|t^{+}\rangle_{i+\rho},$
which gives
$\boldsymbol{S}_{i,1}\cdot\boldsymbol{S}_{i+\rho,0}|t^{+}_{l}\rangle=\textnormal{(non-STD states)}.$
Using the above results, $H^{\prime}|t^{+}_{l}\rangle$ can be evaluated as follows:
$H^{\prime}|t^{+}_{l}\rangle=-\frac{1}{4}\sum_{\rho}J_{\rho}(|t^{+}_{l+\rho}\rangle+|t^{+}_{l-\rho}\rangle)+\textnormal{(non-STD states)}.$
Let $H_{\mathrm{eff}}$ be the effective Hamiltonian of the STD states, in which the energy is measured from the ground-state energy
$E_{\mathrm{g}}=-\frac{3}{4}J_{1}N+\textnormal{(second-order or higher terms)}.$
Then, the matrix elements of $H_{\mathrm{eff}}$ within the first-order perturbation for each model are found to be
$\langle t^{+}_{l^{\prime}}|H_{\mathrm{eff}}^{123}|t^{+}_{l}\rangle\simeq{}$ $\langle t^{+}_{l^{\prime}}|H_{0}|t^{+}_{l}\rangle+\langle t^{+}_{l^{\prime}}|H^{\prime}_{123}|t^{+}_{l}\rangle-E_{\mathrm{g}}\delta_{l^{\prime},l}$
$={}$ $J_{1}\delta_{l^{\prime},l}-\frac{1}{4}\sum_{\rho}J_{\rho}(\delta_{l^{\prime},l+\rho}+\delta_{l^{\prime},l-\rho})$
$={}$ $J_{1}\delta_{l^{\prime},l}-\frac{J_{2}}{4}(\delta_{l^{\prime},l+\boldsymbol{a}_{1}}+\delta_{l^{\prime},l-\boldsymbol{a}_{1}})$
$-\frac{J_{3}}{4}(\delta_{l^{\prime},l+\boldsymbol{a}_{2}}+$ ${}\delta_{l^{\prime},l-\boldsymbol{a}_{2}}+\delta_{l^{\prime},l+\boldsymbol{a}_{1}-\boldsymbol{a}_{2}}+\delta_{l^{\prime},l-\boldsymbol{a}_{1}+\boldsymbol{a}_{2}}),$
$\langle t^{+}_{l^{\prime}}|H_{\mathrm{eff}}^{124}|t^{+}_{l}\rangle\simeq{}$ $\langle t^{+}_{l^{\prime}}|H_{0}|t^{+}_{l}\rangle+\langle t^{+}_{l^{\prime}}|H^{\prime}_{124}|t^{+}_{l}\rangle-E_{\mathrm{g}}\delta_{l^{\prime},l}$
$={}$ $J_{1}\delta_{l^{\prime},l}-\frac{1}{4}\sum_{\rho}J_{\rho}(\delta_{l^{\prime},l+\rho}+\delta_{l^{\prime},l-\rho})$
$={}$ $J_{1}\delta_{l^{\prime},l}-\frac{J_{2}}{4}(\delta_{l^{\prime},l+\boldsymbol{a}_{1}}+\delta_{l^{\prime},l-\boldsymbol{a}_{1}})$
$-\frac{J_{4}}{4}(\delta_{l^{\prime},l+\boldsymbol{a}_{1}-\boldsymbol{a}_{3}}+\delta_{l^{\prime},l-\boldsymbol{a}_{1}+\boldsymbol{a}_{3}}).$
To diagonalize $H_{\mathrm{eff}}$, we introduce a Fourier transform:
$|t^{+}_{\boldsymbol{Q}}\rangle=\frac{1}{\sqrt{N}}\sum_{i}|t^{+}_{i}\rangle e^{-i\boldsymbol{Q}\cdot\boldsymbol{r}_{i}},$
where $N$ is the total number of dimers, and $\boldsymbol{r}_{i}$ is the position vector of the $i$-th dimer. This transformation diagonalizes the effective Hamiltonian and gives the triplet excitation spectrum $\Delta E_{\boldsymbol{Q}}\equiv\langle t^{+}_{\boldsymbol{Q}}|H_{\mathrm{eff}}|t^{+}_{\boldsymbol{Q}}\rangle$:
$\Delta E_{\boldsymbol{Q}}^{123}=J_{1}{}$ $-\frac{J_{2}}{2}\cos(\boldsymbol{Q}\cdot\boldsymbol{a}_{1})$
$-J_{3}\cos\left(\frac{\boldsymbol{Q}\cdot\boldsymbol{a}_{1}}{2}\right)\cos\left[\boldsymbol{Q}\cdot\left(\boldsymbol{a}_{2}-\frac{1}{2}\boldsymbol{a}_{1}\right)\right].$ (11)
Similarly, for the $J_{1}$-$J_{2}$-$J_{4}$ model,
$\Delta E_{\boldsymbol{Q}}^{124}=J_{1}-\frac{J_{2}}{2}\cos(\boldsymbol{Q}\cdot\boldsymbol{a}_{1})-\frac{J_{4}}{2}\cos[\boldsymbol{Q}\cdot(\boldsymbol{a}_{3}-\boldsymbol{a}_{1})].$ (12)
Defining $\boldsymbol{Q}=q_{1}\boldsymbol{b}_{1}+q_{2}\boldsymbol{b}_{2}+q_{3}\boldsymbol{b}_{3}$, we then have
$\boldsymbol{Q}\cdot\boldsymbol{a}_{j}=\sum_{i}q_{i}(\boldsymbol{b}_{i}\cdot\boldsymbol{a}_{j})=\sum_{i}q_{i}(2\pi\delta_{ij})=2\pi q_{j}.$ (13)
Eqs. (11) and (12) become
$\Delta E^{\mathrm{123}}_{\boldsymbol{Q}}={}$ $J_{1}-\frac{J_{2}}{2}\cos(2\pi q_{1})-J_{3}\cos(\pi q_{1})\cos[\pi(2q_{2}-q_{1})],$ (14)
$\Delta E^{\mathrm{124}}_{\boldsymbol{Q}}={}$ $J_{1}-\frac{J_{2}}{2}\cos(2\pi q_{1})-\frac{J_{4}}{2}\cos[2\pi(q_{3}-q_{1})].$ (15)
Figure 8: (a)-(d) Dispersions along $q_{1}$ and $q_{2}$ directions for the $J_{1}$-$J_{2}$-$J_{3}$ model calculated with $J_{1}=16.95$ meV, $J_{2}=-3.9$ meV, and $J_{3}=0.9$ meV. (e)-(h) Dispersions along $q_{1}$ and $q_{3}$ directions for the $J_{1}$-$J_{2}$-$J_{4}$ model calculated with $J_{1}=16.5$ meV, $J_{2}=-6$ meV, and $J_{4}=3$ meV.
A.2 Dynamical structure factor
We denote the ground state by $|\Psi_{\mathrm{g}}\rangle$ and its energy eigenvalue by $E_{\mathrm{g}}$. We also denote the basis of the subspace of wavevectors $\boldsymbol{Q}$ by $\{|\lambda_{\boldsymbol{Q}}\rangle\}$. Let these be eigenstates of $H$ with energy eigenvalues $E_{\lambda\boldsymbol{Q}}$. Then, the dynamical structure factor $\mathcal{S}^{+-}(\boldsymbol{Q},\omega)$ at $T=0$ can be written as
$\mathcal{S}^{+-}(\boldsymbol{Q},\omega)=\sum_{\lambda_{\boldsymbol{Q}}}|\langle\lambda_{\boldsymbol{Q}}|S^{+}_{\boldsymbol{Q}}|\Psi_{\mathrm{g}}\rangle|^{2}\delta(\omega-E_{\lambda_{\boldsymbol{Q}}}+E_{\mathrm{g}}).$
The spin operator $S^{+}_{\boldsymbol{Q}}$ on the right-hand side is defined by
$S^{+}_{\boldsymbol{Q}}=\frac{1}{\sqrt{2N}}\sum_{i,\nu}e^{i\boldsymbol{Q}\cdot\left(\boldsymbol{r}_{i}+\boldsymbol{d}_{\nu}\right)}S^{+}_{i,\nu}$,
where $2N$ represents the total number of spins, $\boldsymbol{r}_{i}$ represents the position of the $i$-th dimer, and $\boldsymbol{r}_{i}+\boldsymbol{d}_{\nu}$ represents the position of the spin $\boldsymbol{S}_{i,\nu}$.
Let the ground state $|\Psi_{\mathrm{g}}\rangle=|s\rangle$. Then, we need to evaluate
$S^{+}_{\boldsymbol{Q}}|s\rangle=\frac{1}{\sqrt{2N}}\sum_{i,\nu}e^{i\boldsymbol{Q}\cdot\left(\boldsymbol{r}_{i}+\boldsymbol{d}_{\nu}\right)}S^{+}_{i,\nu}|s\rangle.$ (16)
Noting that
$S^{+}_{i,\nu}|s\rangle={}$ $S^{+}_{i,\nu}|s\rangle_{i}\prod_{j(\neq i)}|s\rangle_{j}$
$={}$ $(-1)^{\nu+1}|t^{+}\rangle_{i}\prod_{j(\neq i)}|s\rangle_{j}$
$={}$ $(-1)^{\nu+1}|t^{+}_{i}\rangle$,
we find Eq.(16) can be written as
$S^{+}_{\boldsymbol{Q}}|s\rangle={}$ $\frac{1}{\sqrt{2N}}\sum_{i,\nu}e^{i\boldsymbol{Q}\cdot\left(\boldsymbol{r}_{i}+\boldsymbol{d}_{\nu}\right)}(-1)^{\nu+1}|t^{+}_{i}\rangle$
$={}$ $\frac{1}{\sqrt{2}}\sum_{\nu}e^{i\boldsymbol{Q}\cdot\boldsymbol{d}_{\nu}}(-1)^{\nu+1}|t^{+}_{\boldsymbol{Q}}\rangle$, (17)
where
$|t^{+}_{\boldsymbol{Q}}\rangle=\frac{1}{\sqrt{N}}\sum_{i}|t^{+}_{i}\rangle e^{i\boldsymbol{Q}\cdot\boldsymbol{r}_{i}}.$
From Eq. (17) we can see that matrix elements appear only when $|\lambda_{\boldsymbol{Q}}\rangle=|t^{+}_{\boldsymbol{Q}}\rangle$. Therefore, $\mathcal{S}^{+-}(\boldsymbol{Q},\omega)$ can be written as
$\mathcal{S}^{+-}(\boldsymbol{Q},\omega)=|\langle t^{+}_{\boldsymbol{Q}}|S^{+}_{\boldsymbol{Q}}|s\rangle|^{2}\delta(\omega-\Delta E_{\boldsymbol{Q}}).$
The matrix element $\langle t^{+}_{\boldsymbol{Q}}|S^{+}_{\boldsymbol{Q}}|s\rangle$ is
$\langle t^{+}_{\boldsymbol{Q}}|S^{+}_{\boldsymbol{Q}}|s\rangle=\frac{1}{\sqrt{2}}\sum_{\nu}e^{i\boldsymbol{Q}\cdot\boldsymbol{d}_{\nu}}(-1)^{\nu+1}$
and taking the square of the absolute value,
$|\langle t^{+}_{\boldsymbol{Q}}|S^{+}_{\boldsymbol{Q}}|s\rangle|^{2}={}$ $\left|\frac{1}{\sqrt{2}}\sum_{\nu}e^{i\boldsymbol{Q}\cdot\boldsymbol{d}_{\nu}}(-1)^{\nu+1}\right|^{2}$
$={}$ $\frac{1}{2}\left(e^{i\boldsymbol{Q}\cdot\boldsymbol{d}_{0}}-e^{i\boldsymbol{Q}\cdot\boldsymbol{d}_{1}}\right)\left(e^{-i\boldsymbol{Q}\cdot\boldsymbol{d}_{0}}-e^{-i\boldsymbol{Q}\cdot\boldsymbol{d}_{1}}\right)$
$={}$ $1-\cos[\boldsymbol{Q}\cdot(\boldsymbol{d}_{1}-\boldsymbol{d}_{0})].$
Noting that $\boldsymbol{d}_{1}-\boldsymbol{d}_{0}=\frac{2}{3}\boldsymbol{a}_{1}$ for both models, we thus have
$\boldsymbol{Q}\cdot(\boldsymbol{d}_{1}-\boldsymbol{d}_{0})=\frac{4\pi}{3}q_{1},$
and
$|\langle t^{+}_{\boldsymbol{Q}}|S^{+}_{\boldsymbol{Q}}|s\rangle|^{2}=1-\cos\left(\frac{4\pi}{3}q_{1}\right).$
Therefore, the expression for the dynamical structure factor can be written as
$\mathcal{S}^{+-}(\boldsymbol{Q},\omega)=\left[1-\cos\left(\frac{4\pi}{3}q_{1}\right)\right]\delta(\omega-\Delta E_{\boldsymbol{Q}}).$ (18)